OpenStudy (anonymous):

Susan's father is 3 years more than twice as old as Susan. The product of their ages in years is 560. How old is Susan?

OpenStudy (anonymous):

16,35

OpenStudy (anonymous):

fathers age .. y susen age 2x so= 2x+3=y y*x=560 x=560/y so 1120/y +3= y y2 -3y -1120=0 solving this eqn we get y=35 so we get x =16

OpenStudy (anonymous):

Solution:In a theater,there are 1200 seats.The number of seats in each row is 10 less than the number of rows.How many seats are there in each row?